# R1-Track: Direct Application of MLLMs to Visual Object Tracking via Reinforcement Learning

### v1

@@ Abstract
Visual single object tracking aims to continuously localize and estimate the scale of a target in subsequent video frames, given only its initial state in the first frame. This task has traditionally been framed as a template matching problem, evolving through major phases including correlation filters, two-stream networks, and one-stream networks with significant progress achieved. However, these methods typically require explicit classification and regression modeling, depend on supervised training with large-scale datasets, and are limited to the single task of tracking, lacking flexibility. In recent years, multi-modal large language models (MLLMs) have advanced rapidly. Open-source models like Qwen2.5-VL, a flagship MLLMs with strong foundational capabilities, demonstrate excellent performance in grounding tasks. This has spurred interest in applying such models directly to visual tracking. However, experiments reveal that Qwen2.5-VL struggles with template matching between image pairs (i.e., tracking tasks). Inspired by deepseek-R1, we fine-tuned Qwen2.5-VL using the group relative policy optimization (GRPO) reinforcement learning method on a small-scale dataset with a rule-based reward function. The resulting model, R1-Track, achieved notable performance on the GOT-10k benchmark. R1-Track supports flexible initialization via bounding boxes or text descriptions while retaining most of the original model’s general capabilities. And we further discuss potential improvements for R1-Track. This rough technical report summarizes our findings as of May 2025.
@@ 1 Introduction
Since the release of GPT-3 [6], large language models (LLMs) have attracted significant attention and experienced rapid development. Alongside this progress, multimodal large language models (MLLMs) have also advanced considerably, leading to the emergence of high-performing open-source models such as LLaVA [7], InternVL [8], and Qwen2.5-VL [9]. These models are generally capable of processing multiple modalities—including text, images, and videos—and can perform a wide range of downstream tasks effectively, such as object detection, OCR, math and so on.
@@
Unfortunately, our experiments show that Qwen2.5-VL is not effective for target tracking. Specifically, the native Qwen2.5-VL model fails to re-locate a specified target across image pairs, even after extensive prompt engineering attempts.
@@
A straightforward approach is supervised fine-tuning (SFT) of MLLMs, though this may not be optimal. Firstly, SFT risks catastrophic forgetting, which weakens the model’s general capabilities. Secondly, SFT requires hard supervision for every coordinate point (e.g., $[x_{min},y_{min},x_{max},y_{max}]$), making it less suitable for fine-grained grounding tasks. Recent advancements in the application of reinforcement learning (RL) in the domain of LLMs have demonstrated remarkable progress. As evidenced by O1 [10], the implementation of test-time scaling strategies has shown substantial potential in enhancing LLMs’ capacity for complex reasoning. Subsequently, DeepSeek-R1-Zero [11] revealed that even without extensive SFT, the strategic application of a rule-based reward system for reward modeling could effectively harness RL to unlock exceptional reasoning and cognitive capabilities in LLMs.
Accordingly, we also adopt a RL framework to fine-tune MLLMs (MLLMs), employing the GIoU metric as a soft reward function to guide parameter updates. We explore both “think” and “no-think” training strategies. Aligned with Deepseek-R1, we implement group relative policy optimization (GRPO) [12] as the RL optimization algorithm, naming the resulting model R1-Track.
@@
  • We investigate the direct application of MLLMs to visual object tracking tasks, constructing specialized datasets through sampling from the GOT10k training set;
@@
  • We conduct comprehensive experimental analyses, and make our code, datasets, model weights publicly available while discussing future research directions.
@@ 2 Data
Here, we introduce the datasets employed for R1-Track training and testing. For training, we sampled R1-Track-5k, R1-Track-100k and cold-start data. The test data is the GOT-10k [13] official benchmark.
@@
R1-Track-5k is a simple dataset designed for preliminary experimental validation, comprising 5,000 template-search image pairs. We randomly selected 5,000 videos from the GOT-10k training set and extracted two random frames from each video. Centered on the target object in each frame, we performed random cropping followed by resizing the cropped regions to 336×336 pixels. It should be noted that R1-Track-5k has a significant limitation: all target objects in the resized images are forced to a 1:1 aspect ratio.
@@ 3. Return [x_min, y_min, x_max, y_max] coordinates of the target
Cold Start Data.
The SFT and RL fine-tuning processes utilize identical datasets, with differences existing solely in their data assembly formats.
Before conducting SFT or RL, we performed a simple cold start for the base model, which was also achieved through SFT. The cold start dataset consists of 100 “no-think” style data points and only 10 “think” style data points, aiming to equip the model with basic tracking capabilities and accelerate the training process. All subsequent procedures were conducted through further fine-tuning based on the cold-start model. In other words, the training workflow follows the path of “base model >-->- - > cold start >-->- - > SFT/RL".
@@ 3.1 Group Relative Policy Optimization
Group Relative Policy Optimization (GRPO) [12] is a variant of Proximal Policy Optimization (PPO) [14] in reinforcement learning. By comparing groups of candidates responses directly, GRPO eliminates dependency on a critic model and significantly lowers training resources. Given an input question $q$, GRPO first generates $g$ distinct candidate responses $o=\{o_{1},\dots,o_{g}\}$ through policy sampling. The MLLM serves as the reward function to get the corresponding scores $\{r_{1},\dots,r_{g}\}$.
GRPO computes their mean and standard deviation for normalization and determines the quality of these responses:
@@
where $A_{i}$ represents the relative quality of the $i$-th answer. GRPO encourages the model to favor better answers with a high score within the group. The final training objective also considers preventing the optimized policy $\pi_{\theta}$ from deviating far from the base MLLM (step 0) parameters $\pi_{\mathrm{ref}}$ by adding a KL-divergence term $\mathrm{D}_{\mathrm{KL}}$:
@@
The training framework of R1-Track adopts EasyR1 [15], employing a token-level policy gradient Loss similar to DAPO [16].
@@ 3.2 Tracking Rewards
We explore how to use GRPO to enhance the performance of MLLMs in visual object tracking. The overall reward comprises three components: format reward, answer reward, and length reward. And the reward function design varies slightly depending on whether the model incorporates reasoning mechanisms.
@@
To enable the model to output responses in the format we desire. For the “no-think” mode, we expect the model to directly output the standard bounding box coordinates, represented as $[x_{min},x_{max},y_{min},y_{max}]$. In contrast, for the “think” mode, the model should generate outputs in the following format: <think>...</think><answer>$[x_{min},x_{max},y_{min},y_{max}]$</answer>. We use regular expression matching to determine whether the model adheres to our specified format:

### v2

@@ Abstract
Visual single object tracking aims to continuously localize and estimate the scale of a target in subsequent video frames, given only its initial state in the first frame. This task has traditionally been framed as a template matching problem, progressing through correlation filters, two-stream networks, and one-stream networks. However, these methods typically require explicit classification and regression modeling, depend on supervised training on large datasets, and lack flexibility due to their focus on tracking. In recent years, multi-modal large language models (MLLMs) have advanced rapidly. Open-source models such as Qwen2.5-VL, a flagship MLLMs with strong foundational capabilities, demonstrate excellent performance in grounding tasks. This has spurred interest in applying such models directly to visual tracking. However, experiments reveal that Qwen2.5-VL struggles with template matching between image pairs (i.e., tracking tasks). Inspired by DeepSeek-R1, we fine-tuned Qwen2.5-VL using group relative policy optimization (GRPO) reinforcement learning on a small dataset with a rule-based reward function. The resulting model, R1-Track, achieved notable performance on the GOT-10k benchmark. R1-Track supports flexible initialization via bounding boxes or text descriptions while retaining most of the original model’s general capabilities. We further discuss potential improvements for R1-Track. This rough technical report summarizes our findings as of May 2025.
@@ 1 Introduction
Since the release of GPT-3 [6], large language models (LLMs) have attracted significant attention and experienced rapid development. Alongside this progress, multimodal large language models (MLLMs) have also advanced considerably, leading to the emergence of high-performing open-source models such as LLaVA [7], InternVL [8], and Qwen2.5-VL [9]. These models can process multiple modalities, including text, images, and videos, and perform downstream tasks such as object detection, OCR, and mathematical reasoning.
@@
Unfortunately, our experiments show that Qwen2.5-VL is not effective for target tracking. Specifically, even with extensive prompt engineering, the native Qwen2.5-VL model fails to relocate a specified target across image pairs.
@@
Supervised fine-tuning (SFT) of MLLMs is a straightforward but potentially suboptimal approach. First, SFT risks catastrophic forgetting, weakening the model’s general capabilities. Secondly, SFT requires hard supervision for every coordinate point (e.g., $[x_{min},y_{min},x_{max},y_{max}]$), making it less suitable for fine-grained grounding tasks. Recent advancements in applying reinforcement learning (RL) to LLMs have shown remarkable progress. As evidenced by O1 [10], the implementation of test-time scaling strategies has shown substantial potential in enhancing LLMs’ capacity for complex reasoning. Subsequently, DeepSeek-R1-Zero [11] revealed that strategically applying a rule-based reward system for reward modeling can effectively leverage RL to unlock exceptional reasoning and cognitive capabilities in LLMs, even without extensive SFT.
Accordingly, we also adopt a RL framework to fine-tune MLLMs, employing the GIoU metric as a soft reward function to guide parameter updates. We explore both “think” and “no-think” training strategies. Aligned with Deepseek-R1, we implement group relative policy optimization (GRPO) [12] as the RL optimization algorithm, naming the resulting model R1-Track.
@@
  • We investigate the direct application of MLLMs to visual object tracking, constructing specialized datasets by sampling the GOT10k training set;
@@
  • We conduct comprehensive experiments, make our code, datasets, and model weights publicly available, and discuss future research directions.
@@ 2 Data
Here, we introduce the datasets used for R1-Track training and testing. For training, we sampled R1-Track-5k, R1-Track-100k, and cold-start data. The test data is the official GOT-10k benchmark.
@@
R1-Track-5k is a simple dataset designed for preliminary experimental validation, comprising 5,000 template-search image pairs. We randomly selected 5,000 videos from the GOT-10k training set and extracted two random frames from each video. Centering on the target object in each frame, we randomly cropped the images and resized the cropped regions to $336\times 336$ pixels. It should be noted that R1-Track-5k has a significant limitation: all target objects in the resized images are forced to a 1:1 aspect ratio.
@@ 3. Return [x_min, y_min, x_max, y_max] coordinates of the target
Cold-Start Data.
The SFT and RL fine-tuning processes use identical datasets; differences exist solely in their data assembly formats.
Before SFT or RL, we performed a cold start for the base model, also achieved through SFT. The cold-start dataset consists of 100 “no-think” data points and 10 “think” data points to equip the model with basic tracking capabilities and accelerate training. All subsequent procedures involved further fine-tuning of the cold-start model. In other words, the training workflow follows the path of “base model >-->- - > cold start >-->- - > SFT/RL".
@@ 3.1 Group Relative Policy Optimization
Group Relative Policy Optimization (GRPO) [12] is a variant of Proximal Policy Optimization (PPO) [14] in reinforcement learning. By directly comparing groups of candidate responses, GRPO eliminates dependence on a critic model and significantly lowers training resource requirements. Given an input question $q$, GRPO first generates $g$ distinct candidate responses $o=\{o_{1},\dots,o_{g}\}$ through policy sampling. The MLLM serves as the reward function to obtain the corresponding scores $\{r_{1},\dots,r_{g}\}$.
GRPO computes their mean and standard deviation for normalization to determine the quality of these responses:
@@
where $A_{i}$ represents the relative quality of the $i$-th answer. GRPO encourages the model to favor better answers with higher scores within the group. The final training objective also considers preventing the optimized policy $\pi_{\theta}$ from deviating far from the base MLLM (step 0) parameters $\pi_{\mathrm{ref}}$ by adding a KL-divergence term $\mathrm{D}_{\mathrm{KL}}$:
@@
R1-Track’s training framework adopts EasyR1 [15], employing a token-level policy gradient loss similar to DAPO [16].
@@ 3.2 Tracking Rewards
We explore how to use GRPO to enhance the performance of MLLMs in visual object tracking. The overall reward comprises three components: format, answer, and length. Reward function design varies slightly depending on whether the model incorporates reasoning.
@@
To enable the model to output responses in the desired format. In “no-think” mode, we expect the model to directly output standard bounding box coordinates as $[x_{min},x_{max},y_{min},y_{max}]$. In “think” mode, the model should generate outputs in the following format: <think>...</think><answer>$[x_{min},x_{max},y_{min},y_{max}]$</answer>. We use regular expression matching to determine whether the model adheres to the specified format: